-. Let f(x) = 3x + 1 and g(x) = -X- 2. Find (fºg)(4).
Answer #1

* Given that f(x) = 3241 g(x) = -2-2 (fog) (4) f[9(41) f[-4-2] = = f(-6) 31-6) +1 = -18 El - -17 - (fog) (4) = -17-